OpenStudy (anonymous):

integral (0 to x) of sin(t)

OpenStudy (unklerhaukus):

\[\int\limits_0^x\sin(t)\mathrm dt\]

OpenStudy (anonymous):

How do I solve that?

OpenStudy (unklerhaukus):

if you find a standard table of integrals you will see that \[\large\boxed{\int\sin (ax)\mathrm dx=-\frac1a\cos (ax)}\]

OpenStudy (unklerhaukus):

so that means we can do this \[\int\limits_0^x\sin(t)\mathrm dt=-\cos(t)\big|_0^x\]

OpenStudy (anonymous):

=-cos(x)+1

OpenStudy (anonymous):

Thanks

OpenStudy (unklerhaukus):

do you know the next step @mulline ?

OpenStudy (unklerhaukus):

\[f(x)\big|_a^b=f(b)-f(a)\]

OpenStudy (unklerhaukus):

so \[-\cos(t)\big|_0^x=\big(-\cos(x)\big)-\big(-\cos(0)\big)\]

OpenStudy (unklerhaukus):

@agostino direct answers are not helpful

OpenStudy (anonymous):

@UnkleRhaukus One can always check whether his/her answer is correct to see the validity of his/her knowledge.

OpenStudy (anonymous):

Yes I understand how it is done I just making sure I was on the right track

OpenStudy (unklerhaukus):

mulline, you said you didn't know how to solve the question